# More than Carbon: Cradle-to-Grave environmental impacts of GenAI training on the Nvidia A100 GPU

### v1

@@ Abstract
The rapid expansion of artificial intelligence (AI) has intensified concerns about its environmental sustainability. Yet, current assessments predominantly focus on operational carbon emissions using secondary data or estimated values, overlooking environmental impacts in other life cycle stages. This study presents the first comprehensive multi-criteria life cycle assessment (LCA) of AI training, examining 16 environmental impact categories based on detailed primary data collection of the Nvidia A100 SXM 40GB GPU, one of the most widely deployed GPUs for AI training. Our method emphasizes creating primary data through teardown analysis and elemental composition analysis of the individual GPU component groups. To enhance transparency and reproducibility of LCA research on AI, all primary data are made openly accessible. The LCA results for training BLOOM reveal that the use phase dominates 11 out of 16 impact categories including climate change (96%), while manufacturing dominates the remaining 5 impact categories including human toxicity, cancer (99%), eco-toxicity, freshwater (37%) and mineral and metal depletion (85%). For GPT-4’s training process, the use phase dominates 10 out of 16 impact categories, contributing about 96% to both the climate change category and resource use, fossils. The manufacturing stage dominates 6 of 16 impact categories including human toxicity, cancer (94%), eutrophication, freshwater (81%), and resource use, minerals and metals (71%). Assessing the cradle-to-gate environmental impact distribution across GPU component categories reveals that the GPU chip is the largest contributor across 10 out of 16 of impact categories and shows particularly pronounced contributions to climate change (81%), and resource use, fossils (80%). While primary data collection results in modest changes in carbon estimates compared to generic database-derived estimates, substantial variations emerge in other categories. Most notably, abiotic resource depletion of minerals and metals increases by 33%, demonstrating the critical importance of primary data for non-carbon accounting. This multi-criteria analysis expands the Sustainable AI discourse beyond operational carbon emissions, revealing a comprehensive cradle-to-grave ecological footprint that challenges current sustainability narratives and highlighting the need for policy frameworks addressing the full spectrum of AI’s environmental impact.
@@ Introduction
As AI’s existence depends entirely on the availability of powerful computational infrastructure, understanding the training hardware is crucial. Among the various hardware components that enable AI development, GPUs have emerged as the primary workhorses for model training, making their environmental assessment particularly important. Specifically, the release of the Nvidia A100 GPU in 2020 marks a cornerstone of AI training development, becoming one of the most widely deployed GPUs for AI training. The A100 is widely credited with revolutionizing deep learning efficiency [1] by delivering nearly 2.5 times greater computational power and enhanced memory capacity compared to the previous GPU generation. This hardware advancement established a new benchmark for massive AI workloads, enabling the breakthrough model GPT-3 that redefined possibilities in natural language processing [1]. The transformative impact of this technological advancement is reflected in Nvidia’s record-breaking revenues, which showed 125.85% growth from 2023 to 2024 [48, 32] - a direct consequence of unprecedented demand for AI-capable hardware. The availability of such hardware has, in turn, enabled rapid growth in AI model complexity from billions to trillions of parameters driving many of the recent advances in performance. A prominent example is GPT-4, estimated to comprise 1.8 trillion parameters [22], whose scale alone necessitates exponential increases in computational power and consequently greater resource consumption [5, 15]. The combination of growing models and widespread AI adoption has fueled an unprecedented need for cutting-edge data centers equipped with state-of-the-art GPUs, the backbone of modern AI systems. Given the specialized nature and relatively short operational lifespan of these components [36, 14, 21], this fast-paced hardware expansion cycle raises important environmental concerns that extend beyond operational energy consumption.
However, recent environmental assessments predominantly focus on energy consumption during AI model training, with carbon emissions varying based on model size, data center efficiency, and grid carbon intensity [38]. This carbon-centric approach, reinforced by existing literature on the sustainability of AI and regulatory frameworks that prioritize carbon emissions, has intensified the focus on data center energy efficiency while neglecting broader environmental burdens occurring during the manufacturing or disposal of specialized hardware components. Thus, similar to other fields, AI’s environmental assessment is limited by a ‘carbon tunnel vision’ - a narrow focus on carbon emissions that overlooks the wide spectrum of other important environmental and social impacts [23]. Yet, a single-indicator assessment proves insufficient for evaluating the sustainability of AI, as on the one hand, environmental sustainability goes beyond climate change, and on the other, it does not capture burden shifting from one indicator to the other. Currently, most studies employ single-criterion carbon-centric approaches and omit critical life cycle stages such as raw material extraction and the end-of-life. A comprehensive assessment encompassing material extraction, processing, manufacturing, and end-of-life stages is therefore essential for understanding AI’s complete environmental footprint. The few publications offering multi-stage and multi-criteria approaches rely predominantly on secondary data sources and proxy estimates rather than primary data collection, limiting the accuracy and specificity of environmental impacts assessments. Similar data limitations characterize the operational phase, where AI can be either trained or used for inference. While researchers are quantifying AI inference energy consumption, estimates vary significantly based on model size, input length, and prompt complexity [27]. Moreover, declining industry transparency around AI model specifications contributes to the spread of misinformation and contestation of published consumption estimates [28].
@@
Given these pervasive data limitations and transparency challenges across both embodied and operational stages, this study addresses the critical need for comprehensive, openly available primary data to close the fundamental gap in embodied impacts of AI hardware. Therefore, we conduct a comprehensive multi-criteria life cycle assessment (LCA) across all life cycle stages (cradle-to-grave) of the Nvidia A100 SXM GPU based on primary data generated through a two-step methodology. First, we methodically disassembled the GPU into individual component groups through a detailed teardown analysis. Second, we performed a multi-element composition analysis to determine the material composition of each component group. This compositional analysis enables more precise modeling of the different component groups, ultimately leading to more accurate life cycle impacts assessment across the 16 environmental categories. All primary datasets from the teardown and elemental analysis are made openly available to enhance transparency and reproducibility.
@@
This study provides two key contributions to the field of sustainable AI. First, we establish a comprehensive foundation of primary data for AI hardware environmental assessment. This addresses the critical lack of component-specific material inventories that has limited previous studies to secondary data and proxy estimates. Second, we expand the Sustainable AI discourse beyond its narrow focus on operational carbon emissions to include a comprehensive multi-indicator ecological footprint spanning the complete life cycle. Hence, this study offers crucial insights for policymakers developing AI governance frameworks, industry leaders implementing sustainable practices, and researchers advancing environmental assessment methodologies, supporting the development of effective strategies to mitigate AI’s environmental impacts while ensuring sustainable technological progress.
@@
The evolution of AI environmental assessment reflects this diversity in scope and methodology. Early AI sustainability research concentrated exclusively on operational carbon emissions during model training. Strubell et al.(2019) established foundational quantification methods [47], with subsequent studies adopting increasingly sophisticated methodologies. For instance, Lacoste et. al (2019) enhance precision by incorporating geographical energy mix factors for carbon intensity estimation [24], while Dodge et al. (2022) employ location-based and time-specific marginal emissions data per energy unit during model training [7].
@@
Subsequent research expands to include embodied emissions from AI infrastructure manufacturing. Gupta et al. (2021) integrate insights from consumer electronics carbon footprint studies (smartphones, smart speakers, laptops, tablets) to comprehend and address the embodied carbon impact of data center hardware, using industry-reported sustainability data [16]. They suggest that a significant proportion of the carbon footprint is attributable to AI hardware manufacturing [16].
@@
Wu et al. (2022) examine both use phase and manufacturing related carbon footprints across Meta’s industry-scale ML infrastructure using Apple’s 28-core CPU with dual AMD Radeon GPU as a proxy reference due to limited availability of LCA data for actual AI-specific GPUs and accelerators [53]. Their findings demonstrate that embodied carbon represents approximately 50% of the total environmental impact for large-scale ML tasks when renewable energy sources are considered for operational emissions [53]. Li et al. (2025) incorporate embodied carbon considerations for AI infrastructure on the data center-level. The study relies entirely on secondary data sources and proxy estimates from industry reports, most notably using the Dell R740 LCA as a universal baseline and scaling it across different AI system configurations through area-based and power-based scaling factors [25] - a significant methodological limitation compared to studies with direct component-specific LCA data. Their assessment covers cradle-to-gate and the use phase for a single-indicator expressed in carbon emissions. In line with findings from [16] and [53], Li et al. (2025) reveal that manufacturing host processing systems (servers, memory, storage) dominate embodied carbon contributions while GPUs dominate operational carbon, suggesting that optimizing only GPU efficiency misses significant opportunities for carbon reduction [25]. The studies conducted by Gupta et al. (2021), Wu et al. (2022), and Li et al. (2025) highlight the necessity to account for additional life cycle stages beyond the operational phase. Nonetheless, the three studies under consideration are dependent on secondary and proxy data for the embodied emissions calculation. This underscores the critical necessity for the availability of AI-specific data for LCA studies.
@@
Strubell et al. (2019) [47] GPU Card Training NLP model
Lacoste et al. (2019) [24] GPU Card Training ML model
Dodge et al. (2022) [7] GPU Card single AI cloud instance
Gupta et al. (2021)[16] Server not defined secondary & proxy data
Wu et al. (2022) [53] Server Training ML model secondary & proxy data
Li et al (2025) [25] Data Center per AI inference task secondary & proxy data
De Vries et al. (2025) [50] GPU Chip not applicable secondary data
NVIDIA (2025) [34] Server not applicable primary data
Luccioni et al. (2023) [26] Data Center Training AI model secondary & proxy data
Morand et al. (2024) [31] Server AI /ML training instance secondary & proxy data 3
Desroches et al. (2025) [6] IT room Training AI model & AI inference secondary data 5
Schneider et al. (2025) [46] Server Training AI model & AI inference primary data
@@
De Vries (2025) specifically examines GPU chip manufacturing emissions using secondary data from industry reports, manufacturing specifications and market analysis reports to estimate production volumes and technical specifications [50]. However, the study’s reliance on secondary sources and single-indicator carbon assessment limits accuracy compared to primary data collection methods from teardown and elemental analysis. The Nvidia (2025) product carbon footprint report [34] is the first industry-reported cradle-to-gate assessment for an AI server using a combination of primary supplier data of critical components including material composition and production energy consumption data [34]. However, limited transparency in terms of hardware specifications, modeling assumptions, and calculation methodologies, restricts reproducibility and independent verification. The cradle-to-gate plus operational phase approach by Luccioni et al. (2023) establishes precedent for comprehensive AI model assessment [26]. The authors assess BLOOM’s training carbon footprint focusing on the data center-level entirely relying on secondary and proxy data sources from a comparable HPE ProLiant server and estimate GPU embodied emissions due to limited manufacturer disclosure [26].
Morand et al. (2024) model a cradle-to-gate plus operational phase and present the first multi-criteria LCA expressing impacts results in climate change, resource depletion, and primary energy demand [31]. Yet, their assessment leverages only secondary and proxy data, particularly building on existing life cycle inventories, due to the limited availability of primary data for specialized hardware like GPUs. Desroches et al. (2025) follow a cradle-to-gate plus use-phase approach at the IT room level [6]. While this multi-criteria study evaluates five environmental indicator categories - climate change, water use, abiotic resource depletion, primary energy consumption, and final energy consumption - it relies on secondary and proxy data sources [6]. Finally, Schneider et al. (2025) conduct the first comprehensive cradle-to-grave LCA of AI hardware, focusing on Google’s Tensor Processing Units (TPUs) and their complete computing infrastructure at the data center level [46]. The study employs high-quality primary data, supplemented by supplier-provided data for thermal solutions and established LCA databases for standard components, representing the most comprehensive primary data collection for AI manufacturing emissions published to date [46]. However, the assessment is limited to a single-indicator scope and Google’s TPUs, which limits generalizable to other AI hardware use cases.
@@
This section describes the goal and scope and the modeling principles of this study, following ISO 14040:2006 [17] and ISO 14044:2006 [18] standards, ensuring methodological rigor and transparency. An attributional modeling framework is applied to quantify the environmental burdens and resource consumption across multiple impact categories that are specifically attributable to AI hardware production and use. The selection of this approach over a consequential framework is driven by both methodological and practical considerations. The focus of this study is to understand current environmental impacts of training AI models rather than assessing changes in demand for AI over time or market response to increased AI model development. Moreover, given that the secondary objective of this research is to raise awareness about AI hardware’s environmental impacts among non-LCA practitioners, including policymakers, industry stakeholders, and the broader scientific community, the attributional framework’s accessibility and interpretability is particularly well-suited.
@@
Scope This study examines the multi-criteria environmental impacts from cradle-to-grave of training AI services on a high-performance computing (HPC) cluster providing GPU hours. GPU hours are a common metric to measure AI training duration and costs. From its release in 2020 until late 2023, the A100 was the most widely used hardware inside HPC clusters for AI training, playing a crucial role in the development of large-scale AI models. The Life Cycle Impacts Analysis (LCIA) method for the 16 selected criteria rely on the Product Environmental Footprint (PEF) method based on the EU recommendation 2013/179/EU [8], which aligns with ISO 14040/44 standards. Indicators are set for PEF EF 3.0 v2.0 (midpoint) [9]. The modelling approach employed is illustrated in Figure 1.
@@
Functional unit GenAI models do not always fulfill the same purpose, this also holds for BLOOM and GPT-4111unofficial sources estimate the training ranged from 54,000,000- 60,000,00 GPU hours [22].. While BLOOM is a large language model focused on text processing, GPT-4 is a multimodal model capable of processing both text and images. These functional differences make the definition of one functional unit (FU) challenging. Therefore, this study employs a parametric FU with the aim of capturing the essential training characteristics of both models. The FU is defined as:
@@
Normalization In order to provide further support for the communication objective, the LCA results are normalized against a reference situation to eventually provide LCIA normalized results, enabling comparison across diverse environmental indicators that would otherwise have incompatible units and magnitudes. Among several normalization approaches available, we select the Planetary Boundary (PB)-aware normalization defined by Sala et al (2018) [45] as the most relevant for our study objectives. Contrarily to common reference scenario, the PB-aware reference defines annual values of individual impacts aligned with ecological limits, for each impact category. This normalization employs the widely recognized PB framework to enhance the interpretation of results for non-academic stakeholders, thereby facilitating more effective communication of AI’s environmental significance in a global context. Other available LCIA normalization methods typically rely on historical reference scenarios, while the PB-aware normalization anchors our results to the ongoing debate on current planetary limits. Instead of expressing the impacts as a proportion of the actual or historical impacts per capita, this approach expresses the impacts as the proportion of ecological limits per capita considering planetary limits. To contextualize the LCA results within this framework, impacts are normalized to the world population of 2023 [52], as this is the year that GPT-4 was trained in. This representation allows all impact categories to be normalized on the same scale - one inhabitant per year budget. Hence, the normalized metrics express the environmental impacts of training an AI model as the proportion of the annual ‘environmental budget’ per world capita of 2023. However, it is important to note that this per capita normalization does not indicate whether training an AI model is compatible with the global safe operating space (SOS). The results only express the impacts of training an AI model in the equivalence of PB-aware personal budgets. Determining the absolute sustainability of training AI through absolute environmental sustainability assessment (AESA) would require extra-steps such as downscaling/allocation of carrying capacities and the interpretation of absolute sustainability ratios - for instance, scaling the PBs to our FU and comparing against total global AI training activities - which is far beyond the scope of this study.
@@
Figure 2: Annotated components of the Nvidia SXM A100 40 GB GPU showing which components were the subject of a detailed teardown analysis and which underwent an elemental analysis via ICP-OES (author pictures)
Relevant input parameters such as the weight or the surface for each component are recorded. In particular, the die area is an essential input parameter for electronic component for properly scaling the LCI dataset in the Negaoctet database, as it is provided per cm2 of die for most components. The dimensions of the semiconductor die inside the PoPs and largest integrated circuit are revealed by sinking components in a hot sulfuric acid bath, a technique known as chemical decapping [39]. It selectively removed the external packaging, exposing the die area inside. The die areas are then measured using a caliper, achieving a resolution of less than 0.01mm. In the specific case of PoPs, the internal structure contained several ICs. Such a component complexity is far from being captured in current LCA databases, even the ones with specific modeling of electronic components. Therefore, PoPs require extra assumptions to be modeled with best available proxies (see SM). The chemical decapping procedure is not necessary for the GPU die, as it uses a flip-chip package design where the semiconductor die area is only protected by a metallic shield, which is removed using a soldering gun.
@@
In accordance with the defined scope, the environmental impacts during the use phase are solely dependent on the energy consumption of the GPUs. The power consumption of GPUs is contingent upon their utilization. Because AI training is a computationally highly intensive task, we use the maximum thermal design power (TDP) of the A100 SXM (400 W) as the average power consumption per GPU card [33]. This parameter will be subject to a Sensitivity Analysis (SA). When GPUs are not allocated to a training task, the equipment is estimated to consumes an idle power of 85W [49].
@@
The impacts of electricity consumption depend on the composition of the electricity mix where the data center is located. For each FU configuration, we use the impacts per kWh corresponding to the year and location of the training: 2022 in France for BLOOM [26] and 2023, Iowa, USA for GPT-4 [30]. The impacts for the French electricity grid mix are taken from the CODDE database, while the impacts for the Iowa electricity mix are assessed based on the grid mix provided by the Iowa Utilities Commission [19]. The electrical mix considered during GPU chip manufacturing is Taiwan’s energy grid taken from the CODDE database.

### v2

@@ Abstract
The rapid expansion of AI has intensified concerns about its environmental sustainability. Current assessments focus on operational carbon emissions using secondary data, overlooking impacts in other life cycle stages. This study presents a comprehensive multi-criteria life cycle assessment of AI training, examining 16 environmental impact categories using primary data from the Nvidia A100 GPU. Results for GPT-4 training show the use phase dominates 10 categories, contributing 96% to climate change and fossil fuel depletion. Manufacturing dominates 6 categories, including human toxicity (94%) and freshwater eutrophication (81%). The GPU chip is the largest contributor in 10 categories, particularly climate change (81%) and fossil resource use (80%). While primary data produces modest changes in carbon estimates, substantial variations emerge elsewhere, e.g. minerals and metals depletion increases by 33%. This analysis expands Sustainable AI discourse beyond carbon emissions, challenging current sustainability narratives.
@@ Introduction
As AI’s existence depends entirely on the availability of powerful computational infrastructure, understanding the training hardware is crucial. Among the various hardware components that enable AI development, GPUs have emerged as the primary workhorses for model training, making their environmental assessment particularly important. Specifically, the release of the Nvidia A100 GPU in 2020 marks a cornerstone of AI training development, becoming one of the most widely deployed GPUs for AI training. The A100 is widely credited with revolutionizing deep learning efficiency [MediumA100] by delivering nearly 2.5 times greater computational power and enhanced memory capacity compared to the previous GPU generation. This hardware advancement established a new benchmark for massive AI workloads, enabling the breakthrough model GPT-3 that redefined possibilities in natural language processing [MediumA100]. The transformative impact of this technological advancement is reflected in Nvidia’s record-breaking revenues, which showed 125.85% growth from 2023 to 2024 [NVIDIAfiscal2023, NVIDIAfiscal2024] - a direct consequence of unprecedented demand for AI-capable hardware. The availability of such hardware has, in turn, enabled rapid growth in AI model complexity from billions to trillions of parameters driving many of the recent advances in performance. A prominent example is GPT-4, estimated to comprise 1.8 trillion parameters [GPT4_leaked], whose scale alone necessitates exponential increases in computational power and consequently greater resource consumption [debus2023reportingElectr, guo2025deepseek-R1].
The combination of growing models and widespread AI adoption has fueled an unprecedented need for cutting-edge data centers equipped with state-of-the-art GPUs, the backbone of modern AI systems. Given the specialized nature and relatively short operational lifespan of these components [Titan2020, meta2024llama3, karydopoulos2025heatimpact], this fast-paced hardware expansion cycle raises important environmental concerns that extend beyond operational energy consumption. Nonetheless, current environmental assessments predominantly focus on energy consumption during AI model training, with carbon emissions varying based on model size, data center efficiency, and location-based grid carbon intensity [strubell2019, lacoste2019quantifying, dodge2022measuring]. This carbon-centric approach, reinforced by existing literature and regulatory frameworks that prioritize energy consumption, has intensified attention on data center energy efficiency while overlooking broader environmental impacts arising from the manufacturing and disposal of specialized hardware. Thus, similar to other sectors, environmental assessments of AI remain constrained by a ‘carbon tunnel vision’: a narrow focus on carbon emissions that overlooks the wider spectrum of environmental and social impacts [jan_konietzko_moving_2022]. Yet, a single-criterion, carbon-centric approach that omits critical life cycle stages, such as raw material extraction and end-of-life processing, is insufficient for assessing the sustainability of AI. Environmental sustainability extends beyond climate change and requires consideration of potential burden shifting across different impact categories. A comprehensive assessment encompassing material extraction, processing, manufacturing, and end-of-life stages is therefore essential for understanding AI’s complete environmental footprint.
The few publications offering multi-stage and multi-criteria approaches rely predominantly on secondary data sources and proxy estimates rather than primary data collection, limiting the accuracy and specificity of environmental impact assessments (for a detailed review, see the next Section). Similar data limitations affect the operational phase, where AI can be either trained or used for inference. Although recent research has begun quantifying AI inference-related energy consumption, estimates vary widely with model size, input length, and prompt complexity [luccioni2024inference]. Moreover, declining industry transparency around AI model specifications contributes to the spread of misinformation and contestation of published consumption estimates [luccioni2025misinformation].
@@
Given these pervasive data limitations and transparency challenges across both embodied and operational stages, this study addresses the critical need for comprehensive, openly available primary data to close the fundamental gap in embodied impacts of AI hardware. To capture AI’s complete environmental footprint, we conduct the first comprehensive multi-criteria LCA across all life cycle stages (cradle-to-grave) of the Nvidia A100 SXM GPU based on primary data generated through a two-step methodology. First, we methodically disassembled the GPU into component groups through a detailed teardown analysis. Second, we performed a multi-element composition analysis to determine the material composition of each component group. This compositional analysis enables more precise component-level modeling and improves the accuracy of life cycle impacts assessment across the 16 environmental categories. All primary datasets from the teardown and elemental analysis are openly available to enhance transparency and reproducibility [^1] Find the supplementary materials here: github.com/sophia-falk/more-than-carbon, thereby addressing the lack of accessible, high-quality data that has hindered progress in sustainability assessments in the field.
@@
This study provides two key contributions to the field of sustainable AI. First, it establishes a comprehensive foundation of primary data for AI hardware environmental assessment, addressing the critical lack of component-specific material inventories that has limited previous studies to secondary data and proxy estimates. Second, it expand the Sustainable AI discourse beyond its narrow focus on operational carbon emissions to include a comprehensive multi-indicator ecological footprint spanning the complete life cycle. Together, these contributions provide valuable insights for policymakers developing AI governance frameworks, industry leaders implementing sustainable practices, and researchers advancing environmental assessment methodologies, supporting the development of effective strategies to mitigate AI’s environmental impacts while ensuring sustainable technological progress.
@@
The evolution of AI environmental assessment reflects this diversity in scope and methodology. Early AI sustainability research concentrated exclusively on operational carbon emissions during model training. Strubell et al. (2019) established foundational quantification methods [strubell2019], with subsequent studies adopting increasingly sophisticated methodologies. For instance, Lacoste et. al (2019) enhance precision by incorporating geographical energy mix factors for carbon intensity estimation [lacoste2019quantifying], while Dodge et al. (2022) employ location-based and time-specific marginal emissions data per energy unit during model training [dodge2022measuring].
@@
Subsequent research expands to include embodied emissions from AI infrastructure manufacturing. Gupta et al. (2021) integrate insights from consumer electronics carbon footprint studies (smartphones, smart speakers, laptops, tablets) to comprehend and address the embodied carbon impact of data center hardware, using industry-reported sustainability data. They suggest that a significant proportion of the carbon footprint is attributable to AI hardware manufacturing [gupta2021chasing].
@@
Wu et al. (2022) examine both use phase and manufacturing related carbon footprints across Meta’s industry-scale ML infrastructure using Apple’s 28-core CPU with dual AMD Radeon GPU as a proxy reference due to limited availability of LCA data for actual AI-specific GPUs and accelerators. Their findings demonstrate that embodied carbon represents approximately 50% of the total environmental impact for large-scale ML tasks when renewable energy sources are considered for operational emissions [wu2022sustainable]. Li et al. (2025) incorporate embodied carbon considerations for AI infrastructure on the data center-level. The study relies entirely on secondary data sources and proxy estimates from industry reports, most notably using the Dell R740 LCA as a universal baseline and scaling it across different AI system configurations through area-based and power-based scaling factors [Li2025ecoserve] - a significant methodological limitation compared to studies with direct component-specific LCA data. Their assessment covers cradle-to-gate and the use phase for a single-indicator expressed in carbon emissions. In line with findings from [gupta2021chasing] and [wu2022sustainable], Li et al. (2025) reveal that manufacturing host processing systems (servers, memory, storage) dominate embodied carbon contributions while GPUs dominate operational carbon, suggesting that optimizing only GPU efficiency misses significant opportunities for carbon reduction [Li2025ecoserve]. The studies conducted by Gupta et al. (2021), Wu et al. (2022), and Li et al. (2025) highlight the necessity to account for additional life cycle stages beyond the operational phase. Nonetheless, the three studies under consideration are dependent on secondary and proxy data for the embodied emissions calculation. This underscores the critical necessity for the availability of AI-specific data for LCA studies.
@@
Strubell et al. (2019) [strubell2019] GPU Card Training NLP model
Lacoste et al. (2019) [lacoste2019quantifying] GPU Card Training ML model
Dodge et al. (2022) [dodge2022measuring] GPU Card single AI cloud instance
Gupta et al. (2021)[gupta2021chasing] Server not defined secondary & proxy data
Wu et al. (2022) [wu2022sustainable] Server Training ML model secondary & proxy data
Li et al (2025) [Li2025ecoserve] Data Center per AI inference task secondary & proxy data
De Vries et al. (2025) [greenpeace2025] GPU Chip not applicable secondary data
NVIDIA (2025) [NvidiaH100PCF] Server not applicable primary data
Luccioni et al. (2023) [luccioni2023] Data Center Training AI model secondary & proxy data
Berthelot et al. (2024) [berthelot2024estimating] Server AI /ML training secondary & proxy data 3
Morand et al. (2024) [Morand2024MLCA] Server AI /ML training instance secondary & proxy data 3
Desroches et al. (2025) [desroches2025exploringsustainablescalingai] IT room Training AI model & AI inference secondary data 5
Schneider et al. (2025) [schneider2025lifecycleemissionsaihardware] Server Training AI model & AI inference primary data
@@
De Vries (2025) specifically examines GPU chip manufacturing emissions using secondary data from industry reports, manufacturing specifications and market analysis reports to estimate production volumes and technical specifications [greenpeace2025]. However, the study’s reliance on secondary sources and single-indicator carbon assessment limits accuracy compared to primary data collection methods from teardown and elemental analysis. The Nvidia (2025) product carbon footprint report is the first industry-reported cradle-to-gate assessment for an AI server using a combination of primary supplier data of critical components including material composition and production energy consumption data [NvidiaH100PCF]. However, limited transparency in terms of hardware specifications, modeling assumptions, and calculation methodologies, restricts reproducibility and independent verification. The cradle-to-gate plus operational phase approach by Luccioni et al. (2023) establishes precedent for comprehensive AI model assessment. The authors assess BLOOM’s training carbon footprint focusing on the data center-level entirely relying on secondary and proxy data sources from a comparable HPE ProLiant server and estimate GPU embodied emissions due to limited manufacturer disclosure [luccioni2023].
Berthelot et al. (2024) conduct the first multi-criteria LCA expressing impacts results in climate change, resource depletion, and primary energy demand, followed by Morand et al. (2024). Both studies, however, depend on secondary and proxy life cycle inventories due to the limited availability of primary data for specialized hardware such as GPUs. Desroches et al. (2025) follow a cradle-to-gate plus use-phase approach at the IT room level. While this multi-criteria study evaluates five environmental indicator categories - climate change, water use, abiotic resource depletion, primary energy consumption, and final energy consumption - it relies on secondary and proxy data sources [desroches2025exploringsustainablescalingai]. Finally, Schneider et al. (2025) conduct the first comprehensive cradle-to-grave LCA of AI hardware, focusing on Google’s Tensor Processing Units (TPUs) and their complete computing infrastructure at the data center level. The study employs high-quality primary data, supplemented by supplier-provided data for thermal solutions and established LCA databases for standard components, representing the most comprehensive primary data collection for AI manufacturing emissions published to date [schneider2025lifecycleemissionsaihardware]. However, the assessment is limited to a single-indicator scope and Google’s TPUs, which limits generalizable to other AI hardware use cases.
@@
This section describes the goal and scope and the modeling principles of this study, following ISO 14040:2006 [iso14040:2006] and ISO 14044:2006 [iso14044:2006] standards, ensuring methodological rigor and transparency. An attributional modeling framework is applied to quantify the environmental burdens and resource consumption across multiple impact categories that are specifically attributable to AI hardware production and use. The selection of this approach over a consequential framework is driven by both methodological and practical considerations. The focus of this study is to understand current environmental impacts of training AI models rather than assessing changes in demand for AI over time or market response to increased AI model development. Moreover, given that the secondary objective of this research is to raise awareness about AI hardware’s environmental impacts among non-LCA practitioners, including policymakers, industry stakeholders, and the broader scientific community, the attributional framework’s accessibility and interpretability is particularly well-suited.
@@
Scope This study examines the multi-criteria environmental impacts from cradle-to-grave of training AI services on a high-performance computing (HPC) cluster providing GPU hours. GPU hours are a common metric to measure AI training duration and costs. From its release in 2020 until late 2023, the A100 was the most widely used hardware inside HPC clusters for AI training, playing a crucial role in the development of large-scale AI models. The Life Cycle Impacts Analysis (LCIA) method for the 16 selected criteria relies on the Product Environmental Footprint (PEF) method based on the EU recommendation 2013/179/EU [EUreco2013/179/EU], which aligns with ISO 14040/44 standards. Indicators are set for PEF EF 3.0 v2.0 (midpoint) [EPLCA_DeveloperEF]. The modelling approach employed is illustrated in Figure 1.
@@
Functional unit GenAI models do not always fulfill the same purpose, this also holds for BLOOM and GPT-4222unofficial sources estimate the training ranged from 54,000,000- 60,000,00 GPU hours [GPT4_leaked].. While BLOOM is a large language model focused on text processing, GPT-4 is a multimodal model capable of processing both text and images. These functional differences make the definition of one functional unit (FU) challenging. Therefore, this study employs a parametric FU with the aim of capturing the essential training characteristics of both models. The FU is defined as:
@@
Normalization In order to provide further support for the communication objective, the LCA results are normalized against a reference situation to eventually provide LCIA normalized results, enabling comparison across diverse environmental indicators that would otherwise have incompatible units and magnitudes. Among several normalization approaches available, we select the Planetary Boundary (PB)-aware normalization defined by Sala et al (2018) [sala2018_PB_LC_development] as the most relevant for our study objectives. Contrarily to common reference scenario, the PB-aware reference defines annual values of individual impacts aligned with ecological limits, for each impact category. This normalization employs the widely recognized PB framework to enhance the interpretation of results for non-academic stakeholders, thereby facilitating more effective communication of AI’s environmental significance in a global context. Other available LCIA normalization methods typically rely on historical reference scenarios, while the PB-aware normalization anchors our results to the ongoing debate on current planetary limits. Instead of expressing the impacts as a proportion of the actual or historical impacts per capita, this approach expresses the impacts as the proportion of ecological limits per capita considering planetary limits. To contextualize the LCA results within this framework, impacts are normalized to the world population of 2023 [WorldBank2023Pop], as this is the year that GPT-4 was trained in. This representation allows all impact categories to be normalized on the same scale - one inhabitant per year budget. Hence, the normalized metrics express the environmental impacts of training an AI model as the proportion of the annual ‘environmental budget’ per world capita of 2023. However, it is important to note that this per capita normalization does not indicate whether training an AI model is compatible with the global safe operating space (SOS). The results only express the impacts of training an AI model in the equivalence of PB-aware personal budgets. Determining the absolute sustainability of training AI through absolute environmental sustainability assessment (AESA) would require extra-steps such as down-scaling/allocation of carrying capacities and the interpretation of absolute sustainability ratios - for instance, scaling the PBs to our FU and comparing against total global AI training activities - which is far beyond the scope of this study.
@@
Figure 2: Annotated components of the Nvidia SXM A100 40 GB GPU showing which components were the subject of a detailed teardown analysis and which underwent an elemental analysis via ICP-OES (author pictures).
Relevant input parameters such as the weight or the surface for each component are recorded. In particular, the die area is an essential input parameter for electronic component for properly scaling the LCI dataset in the Negaoctet database, as it is provided per cm2 of die for most components. The dimensions of the semiconductor die inside the PoPs and largest integrated circuit are revealed by sinking components in a hot sulfuric acid bath, a technique known as chemical decapping [pirson2021assessing]. It selectively removed the external packaging, exposing the die area inside. The die areas are then measured using a caliper, achieving a resolution of less than 0.01mm. In the specific case of PoPs, the internal structure contained several ICs. Such a component complexity is far from being captured in current LCA databases, even the ones with specific modeling of electronic components. Therefore, PoPs require extra assumptions to be modeled with best available proxies (see SM). The chemical decapping procedure is not necessary for the GPU die, as it uses a flip-chip package design where the semiconductor die area is only protected by a metallic shield, which is removed using a soldering gun.
@@
In accordance with the defined scope, the environmental impacts during the use phase are solely dependent on the energy consumption of the GPUs. The power consumption of GPUs is contingent upon their utilization. Because AI training is a computationally highly intensive task, we use the maximum thermal design power (TDP) of the A100 SXM (400 W) as the average power consumption per GPU card [NvidiaA100Product]. This parameter will be subject to a Sensitivity Analysis (SA). When GPUs are not allocated to a training task, the equipment is estimated to consumes an idle power of 85W [vamja2025partitioning_idle].
@@
The impacts of electricity consumption depend on the composition of the electricity mix where the data center is located. For each FU configuration, we use the impacts per kWh corresponding to the year and location of the training: 2022 in France for BLOOM [luccioni2023] and 2023, Iowa, USA for GPT-4 [microsoft_Iowa]. The impacts for the French electricity grid mix are taken from the CODDE database, while the impacts for the Iowa electricity mix are assessed based on the grid mix provided by the Iowa Utilities Commission [Iowa2023]. The electrical mix considered during GPU chip manufacturing is Taiwan’s energy grid taken from the CODDE database.